OpenStudy (anonymous):

Please show work What % of 48 is 12?

jhonyy9 (jhonyy9):

48 --- 100 12 --- x --------- x= 12*100/48 = 1200/48 = 25 %

OpenStudy (anonymous):

48x = 12 x = 0.25 25%

OpenStudy (anonymous):

What % of 105 is 63 ? Show work